if c varies inversely as the square of d and d=4 when c=0.2, what is the constant of variation
Answer by stanbon(75887)

if c varies inversely as the square of d and d=4 when c=0.2, what is the constant of variation
------
c = k/d^2
0.2 = k/4^2
k = 16*0.2 = 80
-------------------
Cheers,
Stan H.
===================